# linecolor=black,linewidth=0.5pt

### v1

@@ I Introduction
The upcoming integration of ISAC capabilities into a broad range of mobile radio devices (6G and beyond) offers a promising pathway to enhance situational awareness and communication efficiency across various applications. Leveraging the advantages of multi-static sensing within meshed ISAC networks, as introduced in [1], these multi-sensor ISAC architectures comprise various transceiver nodes that act as illuminators and sensors within a distributed infrastructure to allow detection and tracking of position-related parameters of static and moving objects, as illustrated in Fig. 1. A crucial element of this envisioned integration is the development of channel sounding testbeds that include all type of nodes present in such an multi-sensor ISAC network, facilitating the testing of ISAC algorithms across various communication scenarios. A significant hurdle lies in the design and implementation of the mobile airborne sensor nodes to capture the rapid variations and unique propagation characteristics encountered in UAV flight scenarios, especially given the stringent weight and space constraints present at UAV, as comprehensively discussed in the survey by Mao et al. [2]. Ideally, such an experimental testbed should be constructed using COTS hardware to ensure availability across different academic research laboratories and enable customization to various measurement scenarios.
@@
– radar targets44footnotemark: 4 UAV (hexacopter) car, cyclist, UAV (VTOL, hexacopter)
@@ II Multi-node sounding testbed
Figure 2: Simplified block diagrams of the various node types in a COTS-based channel sounding testbed for multi-sensor ISAC networks. The multi-node architecture is modular with a variable number of distributed synchronized transceiver nodes, each featuring configurable signal parameters and switchable receive and transmit paths, as well as an highly accurate ground truth in terms of positioning for all participating nodes.
@@ II-A2 Medium-weight MIMO sensor node
This sensor node utilizes one USRP X310 with two TwinRX frontends, offering a 4-channel receiver with phase synchronization, which is for instance essential for direction-finding applications, and a high dynamic range, eliminating the need for an additional LNA. Alternatively, other RF frontends such as the UBX daughterboards can be employed in the USRP X310 to support transmit and receive operation if required. This receiver node further can be extended by an additional USRP B205mini-i transmitter to provide a reference signal essential for enhancing the accuracy and reliability of direction-of-arrival (DoA) estimation. The USRP X310 is connected via 10 GbE Ethernet, and the USRP B205mini-i via USB 3.0, to a compact high-performance host PC, specifically the Minisforum MS01. This PC weights only 850 g, while featuring dual 10 GbE ports and NVMe SSDs for baseband signal streaming and storage.
@@ II-B Distributed Synchronization
In channel sounding testbeds, synchronization is a critical aspect that ensures the accuracy and reliability of the measurement data and can be broadly divided into two categories:
@@ II-B1 Real-time synchronization
Real-time synchronization is essential for various measurement applications, that require precise timing to coordinate the transmission and reception of signals, for instance for Tx/Rx switching in ISAC networks and using time division multiple access (TDMA) schemes for multi-static radar sensing. GNSS disciplined Oscillatorss (GNSSDOs) typically provide 1 pulse per second (1 PPS) and 10 MHz reference signals synchronizing the clocks of all RF hardware components within the distributed nodes and enabling straightforward synchronization at almost any location with a clear sky view. As introduced in our previous work [5], the real-time synchronization of the heavy-weight sensor nodes can be easily achieved using high-precision timing references such as the Stanford Research Systems FS740 with internal rubidium time base. Implementing synchronization for the mobile nodes, on the other hand, faces additional challenges due to their use in dynamic scenarios, especially when deployed on airborne objects, as well as limitations imposed by slimmed-down RF hardware. To meet the weight and space constraints, a compact, light-weigth GNSSDO, previosuly characterized for UAV deployment [7], the Jackson Labs GPSDO LC__\__XO OCXO, is chosen as a compromise between a small form-factor and comparatively moderate reference signal stability and due its availablity to the authors. This GNSSDO provides both 1 PPS and 10 MHz signals to the SDRs in the light-weigth and medium-weight sensor node. However, the small SDR USRP B205mini-i has only one reference input (REF IN) port, which can accept either the 10 MHz or 1 PPS reference signal. To address this, a remote-controlled RF switch circuit with a latched relay was designed. This circuit, which can be controlled via the Raspberry  Pi 5 GPIO header or USB, allows software-driven toggling between the 1 PPS and 10 MHz reference signal ensuring both initial absolute time synchronization and ongoing relative time synchronization during the measurement. Minor modifications to the original field programmable gate array (FPGA) image of the USRP B205mini-i were made to prevent resettling after switching the reference signal source. The circuit design and FPGA image modifications required are available from our GitHub account [8].
@@ II-B2 Synchronization accuracy after post-processing
While real-time synchronization is crucial for immediate operations, post-processing synchronization accuracy is essential for using the recorded sounding data. Over time, the reference signal standards experience oscillator drift due to environmental factors, such as temperature changes, mechanical disturbances and GNSS signal impairments. This drift can introduce errors into the recorded data, affecting the accuracy of subsequent analyses and localization algorithms. To compensate for this time-variant drifts, typically post-processing techniques are employed by utilizing a temporary beacon Tx with unobstructed LoS to all RXs, as shown in [5]. These techniques involve estimating the synchronization error in the recorded data using high-resolution parameter estimation frameworks to estimate the propagation delay of the line-of-sight (LoS) component and comparing the estimated delay with the expected delayfrom the ground truth data. A fractional delay is then applied to compensate the drift.
@@ II-C Positioning and Orientation
In measurement networks that incorporate flying nodes, the integration of orientation information in conjunction with precise positioning data is crucial for the accuracy of localization outcomes. The 6 degrees of freedom (DOF), while maneuvering freely in three-dimensional space, can significantly alter antenna positions. Highly-precise positioning data with centimeter-level accuracy can be obtained from low-cost RTK GNSS receivers like the u-blox ZED-F9P, using real-time radio technical commission for maritime services (RTCM) correction data from a SAPOS® system, as shown in [5]. Orientation information, derived from an common commercial UAV navigation INS module combining the ICM42688 gyroscope and accelerometer with the RM3100 magnetometer, complements the positioning data by providing the roll, pitch, and yaw angles of the dynamic nodes. This combined information ensures that the exact antenna position and orientation are known at all times, even during complex flight maneuvers.
@@ II-E Measurement Software
The measurement software is designed to record data for offline signal processing, rather than real-time localization, aligning with the testbed’s focus on enabling detailed analysis and algorithm development. The reliable USRP software is responsible for synchronized operation of all SDRs with a RPC interface that allows operators to remotely mute/unmute the Tx paths and start/stop the recording of complex baseband signal at the Rx, as introduced in [5]. This is enhanced using real-time telemetry for continuous supervision and seamless operation. Separate, reusable Python scripts are used to handle telemetry for the numerous GNSSDOs, RTK GNSS receivers, and parameters such as CPU and SSD usage and system temperature for all participating nodes. These scripts run as auto-started system services, ensuring that telemetry data is continuously collected and transmitted via the message queueing telemetry transport (MQTT) protocol. All this is accessible via responsive web applications, allowing operators to monitor the system from mobile devices through a web browser. Given the system’s complexity and the critical nature of the costly and time-consuming large-scale measurement campaigns with a high amount of moving nodes, this real-time telemetry is essential to ensure continuous monitoring and prevent undetected failures from compromising the results.
@@ III Real-world measurement campaign
[…]
– RTK base station SAPOS®
@@
– OFDM, Newman phase sequence
@@
– airborne VTOL, hexacopter (DJI Matrice)
11footnotemark: 1 Heavy-weight MIMO sensor node with tiltable, directional antennas
@@
Figure 4: Simplified illustration of one exemplary scenario of the real-world measurement campaign with stationary sensor nodes (Rx: $\bullet$, Tx: $\bullet$) and mobile sensor nodes at hovering UAVs (Rx: $\bullet$, Tx: $\bullet$) to detect a VTOL aircraft (target: $\bullet$) in arrival and departure situation in terms of radar sensing. The localization error between the estimated target positions, derived using least-squares with estimated target delays of all receiver nodes and sensor node position data, and the actual target position obtained from the ground truth unit, are depicted in the lower right.
To demonstrate the capabilities of our experimental testbed for multi-static radar sensing in ISAC networks, we conducted a field test to exemplarily show channel sounding data set acquisition in diverse multi-node V2X, A2A, and A2G measurement scenarios for different ground-based and airborne targets, covering an area of approximately $\times×100$ m at a 2-story building. The sensor node setup, antenna configuration, trajectory planning, and signal parameters for the real-world measurement campaign were chosen to provide conditions and signal parameters similar to those present for 5G user equipment (UE) in FR1, and are summarized in Table II. The measurement campaign was designed to capture a diverse range of scenarios, with a focus on generating a publicly accessible dataset for ISAC algorithm development and validation. Fig. 4 shows only a small part of the extensive measurement campaign, providing just a sneak peak of the upcoming data set. This specific scenario included two stationary nodes for infrastructure-based sensing: one at the rooftop of the building with one transmitter and two receivers, as well as one at the ground of the parking lot with two receivers, each equipped with tiltable, directional antennas at tripods for exposed positioning. Additionally, the setup included four mobile sensor nodes: one transmitter and three receivers using light-weight sensor nodes mounted on hovering transport hexacopters with omni-directional antennas. The target was a flying VTOL aircraft in arrival and departure situations. All transceiver constellations were calibrated using back-to-back (B2B) calibration.
@@
    applying a fractional delay shift for drift compensation,
@@ IV Conclusions
This paper introduced a synchronized distributed channel sounding testbed designed for investigating and developing algorithms of multi-sensor ISAC networks. The testbed features both airborne and ground-based sensor nodes, with focus on modularity, flexibility, and highly accurate positioning of all participating nodes. The system’s performance was validated through a real-world measurement campaign, demonstrating its capability to detect radar targets such as VTOL aircrafts in an A2A and A2G measurement scenarios using stationary ground-based nodes and mobile sensor nodes at an UAV swarm. Future work will include the public release of data sets to support the development and evaluation of ISAC algorithms.

### v2

@@ I Introduction
The upcoming integration of ISAC capabilities into a broad range of mobile radio devices (6G and beyond) offers a promising pathway to enhance situational awareness and communication efficiency across various applications. Leveraging the advantages of multi-static sensing within meshed ISAC networks, as introduced in [1], these multi-sensor ISAC architectures comprise various transceiver nodes that act as illuminators and sensors within a distributed infrastructure to allow detection and tracking of position-related parameters of static and moving objects, as illustrated in Fig. 1. A crucial element of this envisioned integration is the development of channel sounding [2] testbeds that include all type of nodes present in such an multi-sensor ISAC network, facilitating the testing of ISAC algorithms across various communication scenarios. A significant hurdle lies in the design and implementation of the mobile airborne sensor nodes to capture the rapid variations and unique propagation characteristics encountered in UAV flight scenarios, especially given the stringent weight and space constraints present at UAV, as comprehensively discussed in the survey by Mao et al. [3]. Ideally, such an experimental testbed should be constructed using COTS hardware to ensure availability across different academic research laboratories and enable customization to various measurement scenarios.
@@
– radar targets44footnotemark: 4 UAV (hexacopter) car, cyclist, UAV (VTOL, quadcopter)
@@ II Multi-node sounding testbed
Figure 2: Simplified block diagrams of the various node types in a COTS-based channel sounding testbed for multi-sensor ISAC networks. The multi-node architecture is modular with a variable number of distributed synchronized transceiver nodes, each featuring configurable signal parameters and switchable receive and transmit paths, as well as a highly accurate ground truth in terms of positioning for all participating nodes.
@@ II-A2 Medium-weight MIMO sensor node
This sensor node utilizes one USRP X310 with two TwinRX frontends, offering a 4-channel receiver with phase synchronization, which is for instance essential for direction-finding applications, and a high dynamic range, eliminating the need for an additional LNA [8]. Alternatively, other RF frontends such as the UBX daughterboards can be employed in the USRP X310 to support transmit and receive operation if required. This receiver node further can be extended by an additional USRP B205mini-i transmitter to provide a reference signal essential for enhancing the accuracy and reliability of direction-of-arrival (DoA) estimation. The USRP X310 is connected via 10 GbE Ethernet, and the USRP B205mini-i via USB 3.0, to a compact high-performance host PC, specifically the Minisforum MS01. This PC weights only 850 g, while featuring dual 10 GbE ports and NVMe SSDs for baseband signal streaming and storage.
@@ II-B Distributed Synchronization
In channel sounding testbeds, synchronization is a critical aspect that ensures the accuracy and reliability of the measurement data and can be broadly divided into two categories: real-time synchronization, that is for instance crucial for switching transceiver nodes in ISAC networks, and post-processing synchronization accuracy, which involves drift compensation in recorded measurement data.
@@ II-B1 Real-time synchronization
Real-time synchronization is essential for various measurement applications, that require precise timing to coordinate the transmission and reception of signals, for instance for Tx/Rx switching in ISAC networks and using time division multiple access (TDMA) schemes for multi-static radar sensing. GNSS disciplined Oscillators (GNSSDOs) typically provide 1 pulse per second (1 PPS) and 10 MHz reference signals synchronizing the clocks of all RF hardware components within the distributed nodes and enabling straightforward synchronization at almost any location with a clear sky view. As introduced in our previous work [6], the real-time synchronization of the heavy-weight sensor nodes can be easily achieved using high-precision timing references such as the Stanford Research Systems FS740 with internal rubidium time base. Implementing synchronization for the mobile nodes, on the other hand, faces additional challenges due to their use in dynamic scenarios, especially when deployed on airborne objects, as well as limitations imposed by slimmed-down RF hardware. To meet the weight and space constraints, a compact, light-weigth GNSSDO, previously characterized for UAV deployment [9], the Jackson Labs GPSDO LC_\_XO OCXO, is chosen as a compromise between a small form-factor and comparatively moderate reference signal stability and due its availability to the authors. This GNSSDO provides both 1 PPS and 10 MHz signals to the SDRs in the light-weight and medium-weight sensor node. However, the small SDR USRP B205mini-i has only one reference input (REF IN) port, which can accept either the 10 MHz or 1 PPS reference signal. To address this, a remote-controlled RF switch circuit with a latched relay was designed. This circuit, which can be controlled via the Raspberry  Pi 5 GPIO header or USB, allows software-driven toggling between the 1 PPS and 10 MHz reference signal ensuring both initial absolute time synchronization and ongoing relative time synchronization during the measurement. Minor modifications to the original field programmable gate array (FPGA) image of the USRP B205mini-i were made to prevent resettling after switching the reference signal source. The circuit design and FPGA image modifications required are available from our GitHub account [10].
Figure 3: Time-variant synchronization error illustrated by the LoS delay estimation between one Tx (light-weight sensor node mounted at a hovering UAV) and all Rx nodes (stationary heavy-weight sensor nodes and mobile light-weight sensor nodes at hovering UAVs), as shown in scenario Fig. 5, matching the expected performance range of the used GNSSDOs [6, 9]. High-resolution parameter estimation was applied to estimate the strongest path in the multi-path-affected Rx signals.
@@ II-B2 Synchronization accuracy after post-processing
While real-time synchronization is crucial for immediate operations, post-processing synchronization accuracy is essential for using the recorded sounding data. Over time, the reference signal standards experience oscillator drift due to environmental factors, such as temperature changes, mechanical disturbances and GNSS signal impairments, that can differ significantly between individual sensor nodes, as shown in Fig. 3. This drift can introduce errors into the recorded data in form of sampling clock and carrier frequency offsets, both affecting the accuracy of subsequent analyses and localization algorithms. To compensate for this time-variant drifts, typically post-processing techniques are employed by utilizing a temporary beacon Tx with unobstructed LoS to all RXs, as shown in [6]. These techniques involve estimating the synchronization error in the recorded data using high-resolution parameter estimation frameworks to estimate the propagation delay of the line-of-sight (LoS) component and comparing the estimated delay with the expected delay from the ground truth data. This approach mitigates phase and timing errors induced by the time-variant carrier frequency and sampling clock offsets, so that the post-processed sounding data models the actual physical channel with a greater degree of accuracy.
@@ II-C Positioning and Orientation
In measurement networks that incorporate flying nodes, the integration of orientation information in conjunction with precise positioning data is crucial for the accuracy of localization outcomes. The 6 degrees of freedom (DOF), while maneuvering freely in three-dimensional space, can significantly alter antenna positions. Highly-precise positioning data with centimeter-level accuracy can be obtained from low-cost RTK GNSS receivers like the u-blox ZED-F9P, using real-time radio technical commission for maritime services (RTCM) correction data from a SAPOS® system, as shown in [6], to provide the necessary spatial resolution. Orientation information, derived from an common commercial UAV navigation INS module combining the ICM42688 gyroscope and accelerometer with the RM3100 magnetometer, complements the positioning data by providing the roll, pitch, and yaw angles of the dynamic nodes. This combined information ensures that the exact antenna position and orientation are known at all times, even during complex flight maneuvers.
@@ II-E Measurement Software
The measurement software is designed to record data for offline signal processing, rather than real-time localization, aligning with the testbed’s focus on enabling detailed analysis and algorithm development. The reliable USRP software is responsible for synchronized operation of all SDRs with a RPC interface that allows operators to remotely mute/unmute the Tx paths and start/stop the recording of complex baseband signal at the Rx, as introduced in [6]. This is enhanced using real-time telemetry for continuous supervision and seamless operation. Separate, reusable Python scripts are used to handle telemetry for the numerous GNSSDOs, RTK GNSS receivers, and parameters such as CPU and SSD usage and system temperature for all participating nodes. These scripts run as auto-started system services, ensuring that telemetry data is continuously collected and transmitted via the message queueing telemetry transport (MQTT) protocol. A central server processes the telemetry data, which is sent in form of a single Rx signal period and software telemetry at an update rate of 1 Hz, into real-time spectrum and impulse response plots, providing operators with visual feedback for RF function supervision. Additionally, an interactive map for GNSS function supervision and a telemetry table for color-coded monitoring of over 200 system parameters are generated. All this is accessible via responsive web applications, allowing operators to monitor the system from mobile devices through a web browser. Given the system’s complexity and the critical nature of the costly and time-consuming large-scale measurement campaigns with a high amount of moving nodes, this real-time telemetry is essential to ensure continuous monitoring and prevent undetected failures from compromising the results.
@@ III Real-world measurement campaign
– 2-story building at parking lot $\sim$$\times100$m
– RTK base station SAPOS®[11]
@@
– OFDM Newman phase sequence [12]
@@
– airborne VTOL, quadcopter (DJI Matrice)
11footnotemark: 1 Heavy-weight MIMO sensor node with tiltable, directional antennas with
   10 dB beamwidth at 40 degree.
@@
Figure 5: Simplified illustration of one exemplary scenario of the real-world measurement campaign with stationary sensor nodes (Rx: $\bullet$, Tx: $\bullet$) and mobile sensor nodes at hovering UAVs (Rx: $\bullet$, Tx: $\bullet$) to detect a VTOL aircraft (target: $\bullet$) in arrival and departure situation in terms of radar sensing.
To demonstrate the capabilities of our experimental testbed for multi-static radar sensing in ISAC networks, we conducted a field test to exemplarily show channel sounding data set acquisition in diverse multi-node V2X, A2A, and A2G measurement scenarios for different ground-based and airborne targets, covering an area of approximately $\times100$ m at a 2-story building. The sensor node setup, antenna configuration, trajectory planning, and signal parameters for the real-world measurement campaign were chosen to provide conditions and signal parameters similar to those present for 5G user equipment (UE) in FR1, and are summarized in Table II. The measurement campaign was designed to capture a diverse range of scenarios, with a focus on generating a publicly accessible dataset [13] for ISAC algorithm development and validation.
Mobile and infrastructure-based sensing of flying objects: Fig. 5 shows only a small part of the extensive measurement campaign, providing just a sneak peak of the upcoming data set. This specific scenario included two stationary nodes for infrastructure-based sensing: one at the rooftop of the building with one transmitter and two receivers, as well as one at the ground of the parking lot with two receivers, each equipped with tiltable, directional antennas at tripods for exposed positioning and additional absorbers to minimize LoS power and therefore, to improve receiver sensitivity and linearity. Additionally, the setup included four mobile sensor nodes: one transmitter and three receivers using light-weight sensor nodes mounted on hovering transport hexacopters with omni-directional antennas. The target was a flying VTOL aircraft in arrival and departure situations, as recently used in a study about of bi-static Micro-Doppler analysis in ISAC frameworks [14]. All transceiver constellations were calibrated using back-to-back (B2B) calibration.
Figure 6: Localization error between the estimated target positions, derived using least-squares with estimated target delays of all receiver nodes and sensor node position data, and the actual target position obtained from the ground truth unit.
@@
    applying a fractional delay shift for drift compensation using high-resolution parameter estimation to compare the delay of the observed LoS delay and expected LoS delay calculated from the ground-truth antenna positions,
@@ IV Conclusions
This paper introduced a synchronized distributed channel sounding testbed designed for investigating and developing algorithms of multi-sensor ISAC networks. The testbed features both airborne and ground-based sensor nodes, with focus on modularity, flexibility, and highly accurate positioning of all participating nodes. The system’s performance was validated through a real-world measurement campaign with a publicly available data set, demonstrating its capability to detect radar targets such as VTOL aircrafts in an A2A and A2G measurement scenarios using stationary ground-based nodes and mobile sensor nodes at an UAV swarm. Future work will include the public release of additional data sets, such as for V2X scenarios in parking lots, to support the development and evaluation of ISAC algorithms. Moreover, we will explore advanced techniques to address the synchronization error compensation to ensure more robust and accurate results.